OpenStudy (anonymous):

simplify the radical expression √45π^5

OpenStudy (anonymous):

\[\sqrt{45π^5}=\sqrt{9 \times5 \times π^4 \timesπ } =3π^2\sqrt{5 \times π } =3π^2\sqrt{5 π } \]

OpenStudy (cggurumanjunath):

3*pi^2*sqrt(5*pi)